# Harmonia: Algorithm-Hardware Co-Design for Memory- and Compute-Efficient BFP-based LLM Inference

### v1

@@ IV-D Tiling-aware Dataflow Design
TABLE II: Accuracy results on multiple benchmarks from LongBench v1 (with context length set to 4K). Higher is Better.
@@
Due to the limited capacity of on-chip SRAM, tiling strategies are widely adopted to efficiently support large-scale matrix multiplication. As illustrated in Fig. 15, consider multiplying a matrix A of size $M\times N$ with a matrix B of size $N\times K$, resulting in an output matrix C of size $M\times K$. When adopting dataflow 1 (column-major output flow) to compute matrix C, the total number of EMA for matrices A and B is $\frac{K}{k}\times(M\times N)+N\times K$. In contrast, using dataflow 2 (row-major output flow) results in the total number is $\frac{M}{m}\times(N\times K)+(M\times N)$. Taking the linear layer as an example, for a given model, $N$ and $K$ are fixed, while $k$ and $m$ are determined by the on-chip resource allocation and thus remain constant. However, in practical LLM workloads, the number of input tokens can vary substantially, resulting in a wide dynamic range of $M$. Consequently, the choice between the two dataflows directly impacts the EMA cost under different workloads. Specifically, as $M$ increases, the number of EMA of dataflow 2 gradually approaches and eventually becomes lower than that of dataflow 1.
Therefore, we design a flexible data generation flow (FDGF) controller that can switch between column-major and row-major dataflows through lightweight configuration. This controller manages the read address order of two data dispatchers from on-chip SRAM and coordinates on/off-chip data exchanges. Specifically, under the column-major dataflow, weights remain in on-chip SRAM until all activations are processed, after which the next weight column is fetched from off-chip memory. Conversely, in the row-major dataflow, activations stay on-chip until computation with all weights is completed before being replaced with new activations. This adaptability enables the architecture to select the most memory-efficient dataflow for a given workload, reducing costly EMA and further improving overall system efficiency.
V Evaluation
@@ V-A Experimental Setup
To further evaluate accuracy in more practical scenarios, we evaluate on LongBench. In this setting, we retain the baseline (a) and include KIVI [46], a SOTA algorithm targeting KV cache quantization to reduce memory footprint. We also conduct an ablation study by disabling the asymmetric bit allocation strategy and the offline-online hybrid outlier smoothing technique detailed in Sec. III. This comprehensive setup enables a rigorous assessment of our algorithm.
@@
The synthesis results show that Harmonia occupies an area of 3.53mm2 with a total power of 542 mW. It achieves a peak energy efficiency of 4534 GOPS/W in M8W4/M8M4 mode, and 2267 GOPS/W in M8M8 mode, demonstrating high efficiency under different configurations.
@@ V-B Model Accuracy Evaluation
For PPL evaluation, we report results under two KV cache precision settings: a conservative 8-bit and an aggressive 4-bit configuration, as shown in Table I. With 8-bit mantissas, our method achieves near-lossless accuracy while reducing KV-cache storage by 43.75%, whereas all baselines still use FP16 for attention computation and KV-cache storage. Further truncating the mantissas to 4 bits causes a moderate accuracy drop but yields a 68.75% reduction in storage, representing a practical trade-off between accuracy and memory efficiency.
We further evaluate five instruction-tuned models ranging from 1B to 13B parameters. To better reflect real-world scenarios, their performance is assessed on the LongBench benchmark suite, as summarized in Table II. The original KIVI method quantizes only the KV-cache while keeping weights and other activations in FP16 format, making a direct comparison unfair. To ensure consistency, we implement an extended variant, KIVI-q, which matches the bit-width settings of both weights and activations to those used in Harmonia.
As shown in Table II, our approach incurs only a 0.3% average accuracy drop across five evaluated models compared with the weight-only quantization baseline, while delivering significant gains in both computation and memory efficiency. Notably, for Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct, it even surpasses the baseline accuracy, demonstrating its advantage. Moreover, compared with KIVI-q, Harmonia achieves a 3.9% average accuracy improvement. The ablation study further shows a 5.8% gain over the naïve version, confirming the effectiveness of our algorithm in mitigating accuracy degradation under aggressive compression.
@@ VI Related Work and Discussion
KV cache optimization. During auto-regressive generation, frequent access to the KV cache causes heavy memory and bandwidth pressure, especially in long-context scenarios. Existing methods, such as KV cache quantization [66], [57], [23], [39], reduce storage but introduce runtime overhead. Harmonia instead represents the KV cache in the BFP format, achieving similar compression with lower hardware cost. Moreover, this approach is orthogonal to other techniques, such as eviction [71], [45], [73], and sliding window [15], [64], allowing further compression when combined.

### v2

@@ V Evaluation
TABLE III: Accuracy results on multiple benchmarks from LongBench v1 (context length = 4K). Higher is better.
@@
    For a fair comparison, KIVI-q leverages KIVI’s KV quantization, employing the same weight and activation bit-widths as Harmonia.
    For a fair comparison, MXINT adopts microscaling-based activation quantization, keeping the equivalent bit-widths consistent with Harmonia.
@@ V-A Experimental Setup
To further evaluate accuracy in more practical scenarios, we evaluate on LongBench. In this setting, we retain the baseline (a) and include KIVI [41], a SOTA algorithm targeting KV cache quantization to reduce memory footprint. We also adopt the MXINT[45] format for activation quantization as an additional comparison. Moreover, we preform an ablation study by disabling the asymmetric bit allocation strategy and the offline-online hybrid outlier smoothing technique detailed in Sec. III. This comprehensive setup enables a rigorous assessment of our algorithm.
@@
As summarized in Table I, the synthesis results show that Harmonia occupies 3.38 mm² of silicon area and consumes 300.60 mW of total power. It achieves a peak energy efficiency of 8176 GOPS/W in M8W4/M8M4 mode and 4088 GOPS/W in M8M8 mode, demonstrating consistently high efficiency across different operating configurations.
@@ V-B Model Accuracy Evaluation
For PPL evaluation, we report results under two KV cache precision settings: a conservative 8-bit and an aggressive 4-bit configuration, as shown in Table II. With 8-bit mantissas, our method achieves near-lossless accuracy while reducing KV-cache storage by 43.75%, whereas all baselines still use FP16 for attention computation and KV-cache storage. Further truncating the mantissas to 4 bits causes a moderate accuracy drop but yields a 68.75% reduction in storage, representing a practical trade-off between accuracy and memory efficiency.
We further evaluate five instruction-tuned models ranging from 1B to 13B parameters. To better reflect real-world scenarios, their performance is assessed on the LongBench benchmark suite, as summarized in Table III. Our approach incurs only a 0.3% average accuracy drop across five evaluated models compared with the weight-only quantization baseline, while delivering significant gains in both computation and memory efficiency. Notably, for Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct, it even surpasses the baseline accuracy, demonstrating its advantage. Moreover, compared with KIVI-q and MXINT, Harmonia achieves average accuracy improvements of 3.9% and 7.7%, respectively. The ablation study further shows a 5.8% gain over the naïve version, confirming the effectiveness of our algorithm in mitigating accuracy degradation under aggressive compression.
@@ VI Related Work and Discussion
KV cache optimization. During auto-regressive generation, frequent access to the KV cache causes heavy memory and bandwidth pressure, especially in long-context scenarios. Existing methods, such as KV cache quantization [57], [49] reduce storage but introduce runtime overhead. Harmonia instead represents the KV cache in the BFP format, achieving similar compression with lower hardware cost. Moreover, this approach is orthogonal to other techniques, such as eviction [62], [40], and sliding window [12], [56], allowing further compression when combined.